# Closure for Microscale Fluctuations in Chaotic Mixing Fronts

### v1

@@ II.2 Proposed microscale closure
Combining this argument with Eq. (31) provides a complete closure for microscale scalar variance
@@
$\sigma_{c}^{2}\approx\frac{2D(\nabla\bar{c})^{2}}{\gamma}\left|Ei(-4\pi s_{B}^{\prime}/s_{i})\right|.$ (34)
Eq. (34) indicates that the variance of scalar fluctuations in the presence of a mean scalar gradient weakly depends on the strength of molecular diffusion, since the exponential integral $Ei(x)\sim-\ln x$ when $x<1$.
Alternatively, the closure can be formulated on the concentration microscale (Eq. (19)) is
$s_{m}^{2}\approx 2s_{B}^{2}|Ei(-4\pi s_{B}^{\prime}/s_{i})|.$ (35)
Hence, Eq. (35) appears as a correction factor to the simple anzatz $s_{m}\sim s_{B}$ (Eq. (20)). $s_{m}$ can be larger or lower than $s_{B}$ depending on the ratio of injection versus dissipation scales $s_{i}/s_{B}^{\prime}$. Numerical inspection of Eq. (19) shows that $s_{m}>s_{B}$ when $s_{i}/s_{B}^{\prime}>22.7$, and reversely. Since this ratio can be interpreted as a root-squared Péclet number ($Pe\equiv(s_{B}^{\prime}/s_{i})^{2}$), the critical microscale where $s_{m}\approx s_{B}$ is found around $Pe\approx 515$.
Precising the value of $s_{i}$ is the final ingredient of the microsscale closure problem. We envision $s_{i}$ as the typical scale that divides the dispersive scales above $s_{v}$ from the deformation scales below $s_{v}$. It can be estimated as the scale at which the characteristic time to disperse $(s_{i}^{2}/D)$ is of the order of the characteristic time to deform $(\gamma^{-1})$ :
$s_{i}\sim\sqrt{D/\gamma}\propto s_{V},$ (36)
where we have used Eqs. (2) and (5). Note that $s_{i}$ has the same form as the Batchelor scale (Eq. (1)), but with the diffusion coefficient replaced by the dispersion coefficient. Note that this scale has previously been reported to be important in scalar decay through periodic domains, with mixing rates controlled either by local fluid deformation or by global dispersion, respectively below and above a critical domain size, estimated between $2$ and $3s_{v}$ [17, 38]. Recall finally that the theoretical prediction for the fluctuation variance (or equivalently the fluctuation microscale) relies on the following assumptions: (i) the flow is $\delta$ correlated in time, (ii) $\gamma$ is constant (advection is chaotic and scale independent) and (iii) $s_{B}\ll s_{i}$.
@@
In the following section, we numerically test the accuracy of this prediction in a periodic, two-dimensional, smooth, and chaotic flow, the random sine flow. Transport is solved for a fixed mean gradient in a periodic domain (Sections  III.4 and III.5) and in the case of a dispersing front in an infinite domain (Section III.6).
@@ III Comparison to direct numerical simulations
III.1 Methods
To numerically test Eq.(34), we first consider advection-diffusion of passive scalar in the presence of a fixed large-scale gradient $\nabla\bar{c}\equiv\bm{g}$. As done for Eq.(8), the scalar concentration may be split into mean field and fluctuations
@@ III.1 Methods
thus leading to a source term equal to $-gu(y)\bm{e}_{x}$.
III.2 Velocity field
We consider a generic smooth chaotic velocity field, the two-dimensional Zeldovich sine flow [26], or random wave flow. The sine flow is a continuous transformation operating on a periodic domain $[0,s_{v}]\times[0,s_{v}]$, which is defined as $\bm{u}=(u,v)$, with
@@ III.2 Velocity field
with $T$ the time period of the flow, $j$ the time period number and $U$ a constant velocity magnitude. $\phi_{i},\psi_{i}\in[0,\pi]$ are random phases that change at each time period. The velocity field is divergence free. When $Ku\to 0$ (Eq. (3)), the flow changes rapidly over time, causing minimal fluid deformations, while for $Ku\to\infty$ the persistence of the flow strongly affects the deformation of the fluid elements, which tend to align in the main shear directions. The limit $Ku\to 0$ is the $\delta$-time correlated flow envisioned in Kraichnan [1], where analytical results about exist (Eq. (27)). Without loss of generality, we set $T=1$ and $s_{v}=1$ such that the unique free parameter of the flow is $U=Ku$, which controls both the stretching and dispersion statistics of the flow and its persistence. Particle dispersion in random sine flow is given in the long time limit by [17]
@@
where $\Delta x$ and $\Delta y$ are the mean squared displacement caused by a single flow period. In the Kraichnan limit, the mean stretching rate is given by [25, 1]
@@
while the variance of stretching rates is equal to the mean, e.g. $\sigma_{\gamma}^{2}=\gamma$. Note that $\gamma\to 0$ as $U\to 0$, e.g. in the $\delta$-correlated flow limit. In the case of persistent flows ($UT/s_{V}\gg 0$), material lines partially align with shear directions and such that stretching is less efficient and $\gamma<{\pi^{2}U^{2}T}/({8s_{v}^{2}})$. In addition, the alignement of all material interfaces on the same directions tends to reduce stretching heterogeneity ($\sigma^{2}_{\lambda}<\lambda$). More details about this persistent regime can be found in Meunier and Villermaux [36]. In most random chaotic flows, the elongation of material lines in sine flows approximately follows a log-normal distribution with parameters $\gamma t$ and $\sigma^{2}_{\gamma}t$ that depend on the amplitude $U$, such that Eq. (22) is valid.
III.3 Numerical scheme and convergence
@@
[…]
@@
Figure 2: a) Time evolution of the scalar variance b) Time convergence of the mean $<\sigma^{2}_{c}>$ over 100 realizations of the random phases for various $U$ at fixed $s_{B}^{\prime}=5.3\cdot 10^{-3}$.
@@
After this transient, the simulated scalar field reached a statistically steady state, previously reported as strange eigenmode [39] or persistent pattern [21]. In fact, the asymptotic state of the forced problem is not an eigenmode itself but rather is given by a combination of Floquet modes (Eq. (12)) excited by the source term. The spatial structure of this asymptotic mixing state appears different in the presence or absence of large-scale forcing, as shown in Figure 3, with large-scale gradients clearly appearing in the forced case. This suggests a different repartition of the modes energy in freely decaying or forced scalar transport. In turn, the persistence of the flow, characterised by the value of $UT/s_{v}$, does not strongly affect the heterogeneity of the concentration field (Fig. 3), although it clearly impacts the alignment of material lines with respect to the (vertical) velocity field. Statistical averages are then calculated over time periods of $40\gamma$ and presented below.
III.4 Spectrum at equilibrium
@@
Figure 3: Concentration $c^{\prime}$ for various amplitude $U$ for forced mixing $g=1$ (top and scalar decay $g=0$ (bottom). In the scalar decay case, concentrations are rescaled by the standard deviation.
@@
Figure 4: Comparison between power density spectrum obtained via numerical simulation (dots) and Eq. (27) (lines) at fixed Péclet number for $U=0.1,0.2$ and $0.8$ (subplots a.1) and a.2)) and at fixed $U=Ku$ and various Péclet numbers (subplots b.1) and b.2)). The columns show the same spectrum in log-log and log-lin scale. We used $\gamma/\chi_{0}=\pi$ to normalise $E_{k}$. Note that we normalized $k$ by $1/s_{B}^{\prime}$ in the right column to evidence the universal exponential cutoff.
@@
We plot the spectrum of $c^{\prime}$ for various flow persistence and Péclet numbers in Figure 4. The numerical spectrum clearly shows the $k^{-1}$ scaling at large wavenumber (Figs. 4a.1 and b.1) and the exponential cutoff at small wavenumbers wavenumber (Figs. 4a.2 and b.2) . We test the accuracy of the Kraichnan model (Eq. (27)) combined with the proposed injection mechanism (Eq.(33)) in predicting the numerical spectrum. Given the theoretical stretching statistics and dispersivity of the sine flow (Eqs. (43) and (42)), and the proposed relationship (Eq. (33)), the scalar variance injection rate is
${\chi_{0}}=\gamma\,\frac{g^{2}}{\pi}.$ (48)
[…]
@@
In the case of persistent flows, the theory accurately captures the spectrum in the small-wavenumber limit but fails to describe the high-wavenumber tails, which significantly depart from the exponential prediction. Flow persistence ($U>0$) results in a steeper cutoff than predicted theoretically (Fig. 4 a.2). This may be attributed to the alignment of the material lines in the main shear direction, which tends to limit the heterogeneity of the exponential stretching and push towards a Gaussian cutoff of the spectrum [7]. We will see later that this is, however, not critical for the prediction of scalar fluctuations.
III.5 Scalar fluctuations at equilibrium
@@
In Fig. 5, we plot the scalar fluctuation pdf as a function of flow persistence and Péclet number. Due to the large-scale forcing, the pdf of fluctuation is stationary in time. The pdf shows a characteristic Gaussian shape, as previously observed in forced mixing cases [23, 22]. The Gaussian shape is preserved in all flow persistence and Péclet numbers. Note also that the pdf of scalar fluctuation without large-scale forcing ($g=0$) have thicker tails than the Gaussian (Fig. 5a), which resemble a power law. The Gaussianity of scalar fluctuations in mixing fronts is an important feature because it allows one to accurately capture the distribution of fluctuations by the second moments only, that is, the scalar variance.
@@
Figure 5: Pdf of $c^{\prime}$ as a function of $U$ (Pe=$10^{5}$) and Péclet ($U=0.4$). The black dotted line correspond to the case without large scale forcing, with the concentration normalized by $\sigma_{c}$. Pdf are ensemble averaged over 100 realisations.
In Fig. 6, we report the value of the scalar variance in several flow and transport conditions and compare it with the theoretical closure proposed herein (Eq. (6)). A best fit of the closure (38) to the numerical data gives
$s_{i}\approx 2.7s_{v},$ (49)
for all flow persistence (Fig. 6). This value is close to the critical domain size that divides globally controlled from locally controlled variance decay rates in sine flows (between 2 and $3s_{v}$  [17, 38]), confirming the validity of the proposed scaling Eq. (36). Note that the value of $s_{i}$ may differ in other flows than the sine flow.
Even if Eq. (27) relies on the assumption of $\delta$-correlated flows, we found that the prediction for scalar variance is excellent even in the case of correlations. This is due to the importance of the universal range $k^{-1}$ in determining the integral (16) when $s_{B}^{\prime}\ll s_{i}$. The divergence between curves at large $s_{B}^{\prime}$ (small $(s_{i}/s_{B}^{\prime})^{2}$) is explained by the increasing role of the cutoff shape in the integral, which departs from exponential at finite persistence time (Fig. 4).
@@
Figure 6: Dependence of the variance of scalar fluctuations $\sigma^{2}_{c}$ upon modified Batchelor scale $s_{B}^{\prime}$ (Eq. (28)) and flow amplitudes ($U$) in the sine flow (colored lines with markers). The theoretical approximation Eq. (31) is plotted in thick dashed line while the exact numerical integration of Eq. (34) is plotted as a thin continuous line. In the sine flow, $\chi_{0}/\gamma=1/\pi$ (Eq. (48)) The injection scale $s_{i}$ is fixed to $s_{i}=2.7s_{v}$, to match the numerical data.
@@
It is striking that even at large Péclet numbers, microscale fluctuations remain limited compared to macroscale scalar gradients. For example, at $Pe=1000$, scalar fluctuations are expected to be less than 10% of the macroscopic gradient. At $Pe=10^{8}$, the fluctuations hardly reach 60% of the macroscopic gradient.
III.6 Evolution of a mixing front

### v2

@@ II.2 Scalar variance in mixing fronts
At the micro-scale, Kraichnan [24] shows that scalar fluctuations transferred from large scales obey a well-defined spectral structure (Eq. (11)). We argue that, in freely dispersing fronts, the transfer of energy from macro- to micro-scales operates at a specific scale $s_{i}$, named the injection scale, which separates dispersive processes (well above the velocity scale) from fluid deformation processes (well below the velocity scale). We define $s_{i}$ as the scale at which the characteristic time to destroy variance by dispersion $t_{D}$ is comparable to the characteristic time to destroy variance by stretching-enhanced mixing $t_{\gamma}$. In that sense, it is exactly the scale that separates the globally controlled decay–long wavelength mixing mechanism from the locally controlled decay–short wavelength mixing mechanism described by Haynes and Vanneste [15] and Tsang et al. [14]. Note also that $s_{i}$ could also be associated to the “coarsening scale”[31], below which scalar fluctuations are quickly erased by stretching-enhanced mixing. While this scale has been investigated n the context of multiscale turbulent flows [19], the precise relation to the scale defined here for smooth single scale flows remains to be established.
We estimate the injection scale $s_{i}$ as follows. The characteristic time for variance decay by dispersion in a (closed) domain of length $s_{i}$ is
$t_{D}^{-1}=2\left(\frac{2\pi}{s_{i}}\right)^{2}D$ (17)
while the characteristic time for variance decay by stretching-enhanced mixing is [32]
$t_{\gamma}^{-1}=\frac{\gamma^{2}}{2\sigma^{2}_{\gamma}}.$ (18)
The first equation can be simply obtained with a Fourier transform of the dispersion equation, while the second equation results from log-normal distribution of elongation  [32, 13, 33, 3]. Using the proportionality between mean and variance of stretching rates (Eq. (12)), Eq. (18) simplifies to $t_{\gamma}^{-1}=d\gamma/4$ in $d$-dimensional weakly persistent flow.
Equating $t_{D}$ and $t_{\gamma}$ sets the injection scale to
$s_{i}=4\pi\sqrt{\frac{D\sigma^{2}_{\gamma}}{\gamma^{2}}}.$ (19)
For weakly persistent flows in $d=2$ dimensions, we expect $s_{i}=4\pi\sqrt{D/\gamma}$ and in $d=3$ dimensions, $s_{i}=4\pi\sqrt{2D/(3\gamma)}$. Note the similarity of these expressions for $s_{i}$ and the Batchelor scale, $s_{B}=\sqrt{\kappa/\gamma}$, with the molecular diffusivity $\kappa$ replaced by the dispersion coefficient $D$.
Having defined $\chi_{0}$ and $s_{i}$, the amount of microscale scalar variance in a freely dispersing front is directly obtained by integration of the power density spectrum, e.g.
$\sigma^{2}_{c^{\prime}}=\frac{1}{\pi}\int_{2\pi/s_{i}}^{\infty}E_{k}(k)\,\text{d}k,$ (20)
where $E_{k}$ is given by Eq. (11). Assuming ${F}_{\gamma/\sigma_{\gamma}^{2}}\approx 1$, the integral is analytical, and gives
where $Ei$ is the exponential integral. Note that when integrating Eq. (8), we have implicitly assumed that the scalar spectrum is well approximated by the Kraichnan model up to the scale $s_{i}$, given by Eq. (19). Using the expected scaling for $\gamma$ and $D$ (Eqs. (9),(6)) in the definition of $s_{i}$ (Eq (19), we expect $s_{i}\sim s_{v}$. Our hypothesis is thus justified by the smoothness of the velocity field up to $s_{v}$ (see Fig. 2). While the scalar spectrum likely deviates from the Kraichnan scale close to $s_{i}$, the comparison of Eq. (21) with numerical simulations indicate that this has a limited impact on scalar variance.
Equations (16), (19) and (21) provide a prediction for the magnitude of scalar variance within mixing fronts, based on four independent physical parameters: the molecular diffusion $\kappa$, the dispersivity $D$, the mean and variance of the stretching rate $\gamma$ and $\sigma^{2}_{\gamma}$. Note the absence of any numerical constant that cannot be estimated independently, in contrast to existing theories that describe the power density spectrum in turbulent flows [10]. In Figure 2, we summarise the definitions of all length scales mentioned in the study.
Figure 2: Sketch of the scalar spectrum of a freely dispersing front, with the dispersive spectrum at macroscale (Eq. (13)) and the Kraichnan spectrum at microscale (Eq. (11)). $s_{d}$ is the dispersion scale, $s_{i}$ is the injection scale, $s_{v}$ is the velocity scale, $s_{B}^{\prime}$ is the modified Batchelor scale and $s_{B}$ is the Batchelor scale. The shaded area determines the variability in concentration fluctuations $\sigma^{2}_{c}$
The closure (21) can be used in the macroscale variance evolution equation (8) as follows. Upon dimensional arguments, the dissipation term can be written as
$2\kappa\overline{(\nabla c^{\prime})^{2}}\sim-\sigma^{2}_{c}/\tau_{m},$ (22)
with $\tau_{m}$ a characteristic decay time of scalar variance given by
$\tau_{m}=\frac{\sigma^{2}_{c}}{2D\left(\nabla\bar{c}\right)^{2}}=\gamma^{-1}\left|Ei(-4\pi s^{\prime}_{B}/s_{i})\right|.$ (23)
In the first equality, we have used the assumption of quasi-equilibrium where the left-hand side of Eq. (8) is negligible and, in the second equality, we have used the closure Eq. (21). Note that the ratio $s_{i}/s_{B}^{\prime}$ can be associated to the square root of a Péclet number defined as
@@
$Pe\equiv\left(\frac{s_{i}}{s_{B}^{\prime}}\right)^{2}=8\pi\frac{\sigma^{2}_{\gamma}}{\gamma}\frac{D}{\kappa},$ (24)
which compares dispersive to diffusive fluxes. It is interesting to note that at high $Pe$ (low $s_{B}^{\prime}/s_{i}$), the exponential integral shows a logarithmic growth $|Ei(1/x)|\sim\ln x$, such that Eq. (23) takes a form similar to the /expression of lamellar mixing time in chaotic flows [17], $\tau_{m}\sim\gamma^{-1}\ln Pe_{0}$, with $Pe_{0}=s_{0}^{2}\gamma/\kappa$ a Péclet number defined on the initial size of lamella $s_{0}$. Eqs. (23), (22) and (8) can be used to predict the spatio-temporal evolution of scalar variance in any mixing front.
@@
In the following section, we numerically test the accuracy of these theoretical predictions in a periodic, two-dimensional, smooth, and chaotic flow–the random sine flow–for which $D$, $\gamma$ and $\sigma^{2}_{\gamma}$ can be estimated independently. Transport is solved for a fixed mean gradient in a periodic domain (Sections  III.3 and III.4) and in the case of a dispersing front in an infinite domain (Section III.5).
@@ III Comparison to direct numerical simulations
To numerically test Eq.(21), we first consider the advection-diffusion of passive scalar in the presence of a fixed large-scale gradient $\nabla\bar{c}\equiv\bm{g}$. As done for Eq.(2), the scalar concentration may be split into mean field and fluctuations
@@
thus leading to a source term equal to $gu(y)\bm{e}_{x}$.
III.1 Velocity field
We consider a generic smooth chaotic velocity field, the two-dimensional Zeldovich sine flow [34], or random wave flow. The sine flow is a continuous transformation operating on a periodic domain $0,s_{v}]\times[0,s_{v}]$, which is defined as $\bm{u}=(u,v)$, with
@@ III.1 Velocity field
with $T$ the time period of the flow, $j$ the time period number and $U$ a constant velocity magnitude. $\phi_{i},\psi_{i}\in[0,\pi]$ are random phases that change at each time period. The velocity field is divergence free. When $Ku=UT/s_{v}\to 0$ , the flow changes rapidly over time, causing minimal fluid deformations, while for $Ku\to\infty$ the persistence of the flow strongly affects the deformation of the fluid elements, which tend to align in the main shear directions. The limit $Ku\to 0$ is the $\delta$-time correlated flow envisioned in Kraichnan [24], where Eq. (11) is valid. Without loss of generality, we set $T=1$ and $s_{v}=1$ such that the unique free parameter of the flow is $Ku\equiv U$, which controls both the stretching and dispersion statistics of the flow and its persistence. Particle dispersion in random sine flow is given in the long time limit by [14]
@@
where $\Delta x$ and $\Delta y$ are the mean squared displacement caused by a single flow period (see Supp. Mat. for details). In the Kraichnan limit, the mean stretching rate is given by [26, 24]
@@
while the variance of stretching rates is equal to the mean since the flow is two-dimensional, e.g. $\sigma_{\gamma}^{2}=\gamma$ (Eq. (12) with $d=2$). Note that $\gamma\to 0$ as $U\to 0$, e.g. in the $\delta$-correlated flow limit. In the case of persistent flows ($UT/s_{V}\gg 0$), material lines partially align with shear directions such that stretching is less efficient and $\gamma<{\pi^{2}U^{2}T}/({8s_{v}^{2}})$. In addition, the alignement of all material interfaces on the same directions tends to reduce stretching heterogeneity ($\sigma^{2}_{\lambda}<\lambda$). More details about stretching rates in the persistent regime can be found in Meunier and Villermaux [13].
III.2 Numerical scheme and convergence
@@
followed by a diffusive step (34). The only approximation made by the numerical scheme is the operator splitting. Thus, the algorithm shows only first-order convergence in $\Delta t$, but with a mean relative error smaller than 0.5% when using a large time step $\Delta t=T/2$. The use of GPU makes it extremely efficient, about 2 seconds per time period $T$ for a $20\,000^{2}$ grid with a Nvidia RTX 6000 ADA, which allows us to fully resolve the scalar transport up to Péclet numbers of $10^{8}$ in reasonable time.
@@
Figure 3: a) Time evolution of the scalar variance b) Time convergence of the mean $\sigma^{2}_{c}>$ over 100 realizations of the random phases for various $U$ at fixed $s_{B}^{\prime}=5.3\cdot 10^{-3}$.
@@
After this transient regime, the simulated scalar field reached a statistically steady state, previously reported as strange eigenmode [35] or persistent pattern [36]. In fact, the asymptotic state of the forced problem is not an eigenmode itself but rather is given by a combination of Floquet modes (see Supp. Mat.) excited by the source term. The spatial structure of this asymptotic mixing state appears different in the presence or absence of large-scale forcing, as shown in Figure 4, with large-scale gradients clearly appearing in the forced case. This suggests a different repartition of the modes energy in freely decaying or forced scalar transport. In turn, the persistence of the flow, characterised by the value of $UT/s_{v}$, does not strongly affect the heterogeneity of the concentration field (Fig. 4), although it clearly impacts the alignment of material lines with respect to the (vertical) velocity field. Statistical averages are then calculated over time periods of $40\gamma$ and presented below.
III.3 Spectrum at equilibrium
@@
Figure 4: Concentration $c^{\prime}$ for various amplitude $U$ for forced mixing $g\equiv\nabla\bar{c}=1$ (top and scalar decay $g=0$ (bottom). In the scalar decay case, concentrations are rescaled by the standard deviation. We used a fixed ratio $\gamma/\kappa=5\cdot 10^{5}$ in all simulations.
@@
Figure 5: Comparison between power density spectrum obtained via numerical simulation (dots) and Eq. (11) (lines) at fixed $U=0.1$ and various diffusivity (subplots b.1) and b.2)) and at fixed diffusivity $\kappa=10^{-5}$ for increasing flow persistence (subplots a.1) and a.2)). The columns show the same spectrum in log-log and log-lin scale. Note that we normalized $k$ by $1/s_{B}^{\prime}$ in the right column to evidence the universal exponential cutoff. We used the theoretical value of $\chi_{0}$ obtained for a freely dispersing front (Eq. (16)) to normalize the spectra. Inset of b.1) shows the apparition of spurious frequencies at large flow persistence due to the periodicity of the sine-flow.
@@
We plot the spectrum of $c^{\prime}$ for weakly persistent flow ($U=0.1$) and various Péclet numbers in Figure 5a. The spectrum clearly shows the $k^{-1}$ scaling at small wavenumber (Figs. 5a.1) and an exponential cutoff at large wavenumbers (Figs. 5a.2). Given the theoretical stretching statistics and dispersivity of the sine flow (Eqs. (31) and (30)), and the proposed relationship (Eq. (16)), the scalar variance injection rate is
$\frac{\chi_{0}}{\gamma}=\frac{(s_{v}\nabla\bar{c})^{2}}{\pi},$ (36)
in the limit of $\delta$-correlated flow. Integrating Eq. (36) in Eq. (11), we obtain a theoretical PSD that accurately describes the numerical data in the whole range of scales and Péclet numbers for weakly persistent flows ($U\leq 0.1$) (Figs. 5 b.1 and b.2). This confirms that Eq. (16) correctly captures the large scale injection mechanism in a freely dispersing front (Eq. (19)). Note that the prediction of $\chi_{0}/\gamma$ for the sine-flow involves no fitting parameter in contrast to models developed for turbulent flows that use the Batchelor constant as a fitting parameter to link effective to total strain rates [10].
@@
In the case of persistent flows, the Kraichnan theory accurately captures the spectrum in the small-wavenumber limit (Fig 5b.1) but fails to describe the high-wavenumber tails (Fig 5b.2). We observe that flow persistence results in a steeper cutoff of energies than the exponential cutoff predicted by Kraichnan theory. Such cutoff may be attributed to the alignment of the material lines in the main shear direction, which tends to limit the spatial heterogeneity of the exponential stretching ($\sigma^{2}_{\gamma}<\gamma$) and push towards a Gaussian cutoff of the spectrum, of Batchelor type [23]. Note that at high flow persistence, the scalar spectrum tends to a limiting spectrum when plotted against $ks_{B}^{\prime}$, which is well fitted by
$E_{k}\approx\frac{\chi_{0}}{\gamma k}\exp(-2ks_{B}^{\prime})\exp(-(ks_{B}^{\prime})^{2})$ (37)
which is in essence Eq. (11) augmented by a Gaussian cutoff, à la Batchelor. Note that for $U\geq 1$, spurious peaks appear in the spectrum at small wavenumbers (see inset of Fig 5b.1). These are caused by the periodicity of the sine flow, which creates an artificial resonance for flow amplitudes larger than the domain size. Such peaks would probably disappear in random flows.
III.4 Scalar fluctuations at equilibrium
@@
In Fig. 6, we plot the scalar fluctuation pdf for various flow persistence and Péclet numbers. Due to the large-scale forcing, the pdf of fluctuation is stationary in time. The pdf shows a characteristic Gaussian shape, as previously observed in forced mixing cases [37, 38]. The Gaussian shape is preserved in all flow persistences and Péclet numbers. Note also that the pdf of scalar fluctuations without large-scale forcing ($g=0$) have thicker power-law tails than the Gaussian distribution (Fig. 6a). The Gaussianity of scalar fluctuations in mixing fronts is an important feature because it allows one to accurately capture the distribution of fluctuations from the second moment only, that is, the scalar variance.
@@
Figure 6: Pdf of $c^{\prime}$ as a function of $U$ (with $\kappa=10^{-5}$) and molecular diffusivity (with $U=0.4$). The black dotted line correspond to the case without large scale forcing, with the concentration normalized by $\sigma_{c}$. Pdf are ensemble averaged over 100 realisations.
In Fig. 7, we compare the value of the simulated scalar variance to the theoretical prediction (Eq. (7)) and its approximation with ${F}_{\gamma/\sigma^{2}_{\gamma}}\approx 1$ (Eq. (21)). The injection scale (Eq. (19)) is estimated using Eq. (30) and Eq. (31) for the dispersion coefficient and the stretching rate respectively, giving
$s_{i}=2\sqrt{2}\,s_{v}\approx 2.8\,s_{v}.$ (38)
Note that this value agrees with the critical domain size (between 2 and $3s_{v}$) that divides globally controlled from locally controlled variance decay rates in 2D sine flows [14, 39]. The predicted variance (Eq. 21) accurately captures the numerical data over a large range of molecular diffusivity values and flow amplitudes. The approximation ${F}_{\gamma/\sigma^{2}_{\gamma}}\approx 1$ is valid even in two-dimensional flows because it is mainly the universal $k^{-1}$ regime at low wavenumbers and the exponential cutoff at high wavenumbers that determines the value of the integral (21), rather than the exact power-law scaling at intermediate wavenumbers. While the Kraichnan theory rely on $\delta$-correlated flows, we find that the prediction for scalar variance is excellent even in the case of correlations for $Pe>10^{3}$. This is due to the importance of the universal range $k^{-1}$ that emerges both in correlated and uncorrelated flow, which determines the integral (8) when $s_{B}^{\prime}\ll s_{i}$. The divergence between curves at large $s_{B}^{\prime}$ (small $Pe$) is explained by the increasing role of the cutoff shape in the integral, which departs from exponential at finite persistence time (see Fig. 5). As shown in Fig. 7, using the modified spectrum (37) in Eq. (8) to account for flow persistence only captures part of the variance drop observed below $Pe=10^{3}$. Thus, we conclude that in addition to a change in the spectrum, flow persistence may also push the injection scale $s_{i}$ to smaller values.
@@
Figure 7: Dependence of the variance of scalar fluctuations $\sigma^{2}_{c}$ upon the Péclet number (Eq. (24)) and flow amplitudes ($U$) in the sine flow (colored lines with markers). The approximation Eq. (21) is plotted in thick dashed line while the exact numerical integration of Kraichnan spectrum (Eq. (8)) is plotted as a thin continuous line. In the sine flow, $\chi_{0}/\gamma=(s_{v}\nabla\bar{c})^{2}/\pi$ (Eq. (36)) and $s_{i}=2\sqrt{2}s_{v}$ (Eq. (38)). The dashed line represents the numerical integration of the modified Kraichnan spectrum (Eq. (37)) observed in the case of persistent flows.
@@
Note that even at large Péclet numbers, microscale fluctuations remain limited compared to macroscale scalar gradients. For example, at $Pe=10^{3}$, scalar fluctuations are expected to be less than 10% of the macroscopic gradient. At $Pe=10^{8}$, the fluctuations hardly reach 60% of the macroscopic gradient.
III.5 Evolution of a mixing front